OpenStudy (anonymous):

A pendulum in an antique clock swings above a tabletop. The number of centimeters, C, that the tip of the pendulum is from the tabletop is a function of time, t, in seconds. The function that models the distance of the tip of the pendulum from the tabletop is C(t) = 2(cos 4πt) + 12. How high is the pendulum when t = 1/4 second.

OpenStudy (anonymous):

Here, simply plug 1/4 in for t. \[\frac{4\pi}{1} \frac{1}{4} = \frac{4\pi}{4} = \pi\]

OpenStudy (anonymous):

replace 1/4 for t?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so whould i just to do order of operations from there?

OpenStudy (anonymous):

Yes, so find what the argument for cosine is.

OpenStudy (anonymous):

i got 14

OpenStudy (anonymous):